OpenStudy (anonymous):

x+5/3x+2 - 4-2x/3x^2+8x+4 = x+4/ x+2

hero (hero):

\[\frac{x+5}{3x+2} - \frac{4 - 2x}{3x^2 + 8x + 4} = \frac{x+4}{x+2}\]

OpenStudy (anonymous):

yes

hero (hero):

Are you sure?

OpenStudy (anonymous):

Absolutely.

hero (hero):

The steps are pretty much the same as last time. You do have to factor the quadratic first though.

OpenStudy (anonymous):

Here's the deal. I need a step by step, detailed solution. I need it so I can use at as an example for all of my other problems. Excuse me for my terse words and irritability but I don't have time to thoroughly do it on my own, seeing as I have more important things to do. It would be much appreciated if someone who could do what I have asked actually does it. Thank you. Don't take this personally.

hero (hero):

I gave you an example for the last one

hero (hero):

And i already told you that it is similar to this.

hero (hero):

You have more important things to do than your homework? Like what? skateboarding?

OpenStudy (anonymous):

Well, Mr. Sass, alright. No big deal. Moving on.

OpenStudy (anonymous):

:)